OpenStudy (anonymous):

Help please...Find the exact value of the trigonometric function. csc 7π/6

OpenStudy (jdoe0001):

check your Unit Circle, keep in mind that \(\bf csc\left(\frac{7\pi}{6}\right) \implies \cfrac{1}{sin\left(\frac{7\pi}{6}\right)}\) so, find \(\bf sin\left(\frac{7\pi}{6}\right)\) and flip it upside-down

OpenStudy (anonymous):

thank you